A manufacturer knows that their items have a normally distributed lifespan, with a mean of 11.1 years, and standard deviation of 3.3 years.
The 6% of items with the shortest lifespan will last less than how many years?
Give your answer to one decimal place.

Answer by Boreal(15235)

z(0.06)=-1.555
-1.555=(x-11.1)/3.3 since z=(x-mean)/sd
-5.13=x-11
x=5.87 years or 5.9 years